A car used 7 gallons of gasoline to go a distance of 210 miles. At that rate, how much gasoline would the car use to go 495 miles?
Found 2 solutions by rfer, Alan3354:
Answer by rfer(16322)

7/120=x/495
120x=3465
x=28.9 gal

Answer by Alan3354(69443)

210/7 = 30 mph
495/30 = 16.5 gallons
-----------
or
210:7::495:x
210x = 7*495
x = 7*495/210 = 16.5 gallons
-----------
The other tutor made a typo, did 120 miles, but that'll be useful if you plan a 120 mile trip sometime.